### v2

# Guided Query Refinement: Multimodal Hybrid Retrieval with Test-Time Optimization

Abstract
Multimodal encoders have pushed the boundaries of visual document retrieval, matching textual query tokens directly to image patches and achieving state-of-the-art performance on public benchmarks. Recent models relying on this paradigm have massively scaled the sizes of their query and document representations, presenting obstacles to deployment and scalability in real-world pipelines. Furthermore, purely vision-centric approaches may be constrained by the inherent modality gap still exhibited by modern vision-language models. In this work, we connect these challenges to the paradigm of hybrid retrieval, investigating whether a lightweight dense text retriever can enhance a stronger vision-centric model. Existing hybrid methods, which rely on coarse-grained fusion of ranks or scores, fail to exploit the rich interactions within each model’s representation space. To address this, we introduce Guided Query Refinement (GQR), a novel test-time optimization method that refines a primary retriever’s query embedding using guidance from a complementary retriever’s scores. Through extensive experiments on visual document retrieval benchmarks, we demonstrate that GQR allows vision-centric models to match the performance of models with significantly larger representations, while being up to 14x faster and requiring 54x less memory. Our findings show that GQR effectively pushes the Pareto frontier for performance and efficiency in multimodal retrieval. We release our code here.
1 Introduction
Visual document retrieval is the task of returning relevant documents – typically PDFs containing figures, tables, and other visual elements – in response to a textual query (mathew2021docvqadatasetvqadocument; mathew2021infographicvqa; li-etal-2024-multimodal-arxiv; zhu2022towards; faysse2025colpali). To tackle this task, neural retrieval pipelines often follow a text-centric approach, relying on OCR or vision-language models to convert source documents into textual chunks, and then constructing an index using semantic text encoders (karpukhin-etal-2020-dense; VisualMRC2021). An alternative, vision-centric, approach relies instead on multimodal encoder models. Building on the ColBERT (khattab2020colbert) late-interaction approach, ColPali-based (faysse2025colpali) encoders operate directly on image patches, and yield multi-vector embedding representations of images and queries.
While this approach achieves state-of-the-art results on public benchmarks of visual document retrieval [^1] https://huggingface.co/blog/manu/vidore-v2 (mace2025vidore), open challenges within this paradigm remain. First, to pursue state-of-the-art performance, recent late-interaction multimodal retrievers [^2] Henceforth, we use “retrievers” and “encoders” interchangeably. For additional background on these emerging retrieval paradigms, refer to Appendix A. massively scale the length and dimensionality of query and document representations. This can incur a substantial latency and storage overhead, hindering the ability to provide an efficient and scalable solution. For example, Llama-NemoRetriever-Colembed-3B represents each document page with 10 MB of memory (xu2025llama_nemo), three orders of magnitude more than single-vector dense retrievers (Table 4). Secondly, a vision-centric approach for matching textual queries to textually rich documents may be limited by the substantial modality gap (clavié2025readbenchmeasuringdensetext; li2025closingmodalitygapmixed; role2025gapquantifyingreducingmodality) exhibited by modern vision-language models. These gaps motivate exploring complementary approaches for improving the performance of multimodal encoders.
An early concept in the application of neural retrievers has been that of hybrid retrieval (dos-santos-etal-2015-learning; kuzi2020leveraging), where the outputs of different retrievers are aggregated at the level of ranks or query-document similarity scores (see Figure 1) to obtain the final list of retrieved documents. Hybrid retrieval (or hybrid search) most commonly refers to the combination of a neural semantic text retriever with a sparse lexical representation (e.g., BM25). More broadly, it reflects the notion that models relying on different types of representations can capture complementary aspects of the data, which can in turn be leveraged to boost the overall performance on the task.
In this work, we seek to connect these two threads, testing whether the paradigm of hybrid retrieval can complement modern multimodal encoders. Dense text retrievers typically have low online latencies and incur a small storage footprint, and they provide a uni-modal signal between the query and the document representation. Thus, hybrid retrieval between text and image encoders emerges as a natural candidate for enhancing the performance of a multimodal retrieval system. However, in this work we argue that standard hybrid retrieval methods rely on a rather coarse-grained view of the perspective of each retriever – they cannot utilize the rich query-document interactions within the model representation space (Figure 1).
Figure 1: Hybrid retrieval methods. Aggregating the outputs of two retrievers is typically done at the level of ranks (§2.1) or scores (§2.2). Utilizing the information of both representations effectively and efficiently is difficult to achieve in practice. Here we propose a novel approach of Guided Query Refinement (GQR), using similarity scores from an complementary retriever (left) at test time, to inform the query representation of a primary retriever (right).
Aiming to harness this untapped potential of hybrid retrieval, in this work we propose Guided Query Refinement (GQR) (Figure 2), a novel approach for aggregating retriever outputs. Given a query at test time, GQR iteratively optimizes the query representations of a primary retriever with gradient descent, using similarity scores from a complementary retriever. The refined query representation softly incorporates the complementary retriever’s signal, remaining subject to the query–document interactions in the primary retriever space. This updated query embedding is then used to score the documents and return an updated document list. Notably, GQR is architecture-agnostic and can be used to optimize both single- and multi-vector query embeddings.
We conduct extensive experiments on established visual retrieval benchmarks, evaluating nine pairs of state-of-the-art vision and text retrievers and comparing GQR to standard hybrid retrieval approaches. Our results (§3.2) demonstrate consistent gains for models using GQR over base models and other hybrid baselines. Despite the fact that text-centric models achieve lower performance on the task, we find that the complementary signal they provide through GQR proves useful for ColPali-based models. On ViDoRe 2, Colnomic-Embed-Multimodal-7B with GQR is nearly on par with Llama-NemoRetriever-Colembed-3B, while being $\approx\!\times 14$ faster and requiring $\approx\!\times 54$ less memory, and outperforms it while being $\approx\!\times 7$ faster and requiring $\approx\!\times 24$ less memory. Our results and analysis establish that ColPali-based methods using GQR are on the latency and memory Pareto-fronts on the task of visual document retrieval.
2 Methods
Hybrid retrieval variants can be organized into three conceptual levels, reflecting the granularity in which test-time aggregation is performed (Figure 1): the level of document rankings, the level of query-document similarity scores, or the level of embedding representations. The earlier the aggregation, the more information is available, and the more informative the exchange between models can be; however, richer information also increases the burden of normalization and geometrical alignment across spaces.
We begin this section by outlining prominent methods at each level (§2.1, §2.2). We explain that while early representation-level aggregation could be desired due to its richness, it is difficult to achieve in practice. We then present our method, Guided Query Refinement (GQR), which lies between the levels of scores and of representations (§2.3) and provide the motivation for it (§2.4).
Notations.
Given a query $q$ and retriever $m$, we denote the representation of $q$ by $m$ as $e^{q}_{m}$. Similarly, given a set of documents $D=\{d_{i}\}_{i=1}^{N}$, we have $e^{d_{i}}_{m}$ for all $d_{i}\in D$. Document relevance to the query is estimated using a similarity score $s_{m}(q,d_{i})$ between the representations $e^{q}_{m}$ and $e^{d_{i}}_{m}$, typically via cosine similarity (Equation 3) or MaxSim (Equation 4).
$\pi_{m}(q)$ denotes the list of the documents returned by retriever $m$ for query $q$. $K$ is the length of the retrieved list of documents, and we assume it is constant across retrievers. $\mathrm{rank}_{m}(d)$ is the 1-indexed position of $\pi_{m}(q)$ after sorting $\pi_{m}(q)$ by the scores $s_{m}(q,\cdot)$ in descending order. If $d\notin\pi_{m}(q)$, then $\mathrm{rank}_{m}(q,d)=K{+}1$.
Finally, while the formulation is general and applies to any number of retrievers $M$, in this work we focus on the case of $M=2$.
2.1 Ranking-Level Aggregation
Ranking-level aggregation is the simplest form of information exchange between retrievers: each query and document pair is reduced to a single integer rank. While limited in its expressivity, it requires no extra normalizations or alignments, and is therefore widely used in production pipelines [^3] Milvus docs; Elasticsearch docs..
Reciprocal Rank Fusion (RRF).
RRF (36196) combines ranked lists by weighting each item based on the reciprocal of its rank. The RRF constant $\kappa>0$ dampens the impact of very high ranks and controls how much credit is given to mid-list occurrences.
$\mathrm{RRF}(d)=\sum_{m=1}^{M}\frac{1}{\,\kappa+\mathrm{rank}_{m}(d)\,}$ (1)
We also consider Average Ranking, which directly averages ranks across retrievers; see Equation 6 for the formal definition.
2.2 Score-Level Aggregation
Score-level aggregation operates one step deeper than ranking aggregation, operating on the real-valued similarity scores $s_{m}(q,d_{i})$ between the query and documents. To ensure that the scores of different retrievers are in the same scale and range, the common practice (bruch2023analysis) is to first apply a normalization function $N_{m}$ – yielding $\tilde{s}_{m}(q,d_{i})$ – and then aggregate across retrievers:
$\tilde{s}_{m}(q,d_{i})=N_{m}\!\left(s_{m}(q,d_{i})\right),\qquad\mathrm{Score}(q,d_{i})=\frac{1}{M}\sum_{m=1}^{M}\tilde{s}_{m}(q,d_{i}).$ (2)
In this work, we evaluate two variants with different normalizations, Score Aggregation (Min-Max) and Score Aggregation (SoftMax). See Appendix C for details.
Tuned Variants.
More generally, the above methods can be viewed as a weighted aggregation, where the examples above are the uniform case, with each retriever assigned a weight $\alpha=1/M$ (for two retrievers, $\alpha=0.5$ for each). Given a development set, these weights can be fit (bruch2023analysis). Here, for $M=2$, the two retrievers are assigned relative weights $\alpha$ and $1-\alpha$, yielding the parameterized variants Average Ranking - Tuned, RRF - Tuned, Score Aggregation (Min-Max) - Tuned, and Score Aggregation (SoftMax) - Tuned. Details are in Appendix C.
2.3 Guided Query Refinement
Representation-level information carries the richest potential for effective aggregation. Embedding-level projections that align representational spaces are used extensively in modern vision–language systems to combine visual and textual inputs (radford2021clip; jia2021align; li2021albef; li2022blip). At test time, however, operating directly on representations is hindered by heterogeneity: encoders may use a single vector or many vectors per document and query, and they operate within differing dimensionalities and scales. Thus, with strict latency and memory budgets and without access to supervision, aggregation at this level is not trivial. [^4] Concatenating single-vector embeddings is feasible, yet under dot-product scoring this reduces to an unnormalized sum of separate scores, and does not enable interaction between the spaces.
Our goal in this work is to exploit the rich information in query and document representations while remaining architecture-agnostic, lightweight, and practical. To this end, we propose Guided Query Refinement (GQR), a novel method for combining the outputs of two retrievers – a primary retriever $m_{1}$ and a complementary retriever $m_{2}$ (Algorithm 1, Figure 2). GQR refines $m_{1}$’s query representation based on the signal of $m_{2}$’s scores.
At inference time, given a user query $q$, an index search is run with each retriever to obtain its top-$K$ document list $\pi_{m}(q)$. The union of these lists, $\mathcal{C}(q)=\bigcup_{m=1}^{M}\pi_{m}(q)$, serves as the candidate pool of documents. For each retriever $m_{j}\in\{m_{1},m_{2}\}$, we define a distribution over $\mathcal{C}(q)$ via a Softmax:
$p_{j}(d_{i}\mid e_{j}^{q})=\frac{\exp\big(s_{j}(q,d_{i})\big)}{\sum_{k=1}^{|\mathcal{C}(q)|}\exp\big(s_{j}(q,d_{k})\big)}\quad\text{for }i=1,\ldots,|\mathcal{C}(q)|.$
We denote the initial query embedding of $m_{1}$ by $z^{(0)}=e_{1}^{q}$ (GQR is applicable to both single and multi-vector embeddings), and we update it at each step $t$, $z^{(t)}$ for $T$ steps. At step $t$, the consensus distribution of $m_{1}$ and $m_{2}$ is defined as
$p_{\mathrm{avg}}^{(t)}(d)=\tfrac{1}{2}\Big(p_{1}\big(d\mid z^{(t)}\big)+p_{2}\big(d\mid e_{2}^{q}\big)\Big),$
such that only $p_{1}$ depends on $t$ through $z^{(t)}$ and $p_{2}$ is fixed by $e_{2}^{q}$.
We minimize
$\mathcal{L}^{(t)}=\mathrm{KL}\!\big(p_{\mathrm{avg}}^{(t)}(d)\,\|\,p_{1}(d\mid z^{(t)})\big).$
Here, $\mathrm{KL}$ is the Kullback–Leibler divergence (Equation 12).
We apply a gradient step on the query representation with step size $\alpha$ [^5] We define GQR with gradient descent for simplicity, but in practice we found Adam (kingma2015adam) to perform better and use it as the optimizer.,
$z^{(t+1)}=z^{(t)}-\alpha\,\nabla_{z}\mathcal{L}\!\big(z^{(t)}\big),$
where $T$ and $\alpha$ are hyperparameters.
We then compute the final scores from retriever $m_{1}$,
$s_{1}^{(T)}(q,d)=s_{1}\!\big(z^{(T)},d\big)\quad\text{for }d\in\mathcal{C}(q).$
Finally, we produce the list of retrieval results by sorting $\mathcal{C}(q)$ in descending order of $s_{1}^{(T)}(q,d)$, returning the first $K$ elements.
Figure 2: Guided Query Refinement (GQR). Stage 1: Two retrievers independently encode the query and retrieve top-$K$ documents, forming a candidate pool. Stage 2: The primary query embedding is iteratively refined ($z^{(t)}$) over $T$ iterations, by minimizing the KL divergence between a consensus distribution and the primary distribution.
2.4 GQR - Motivation and Rationale
Our approach is inspired by test time optimization methods that rely on pseudo-relevance feedback from a stronger cross-encoder (yu2021improvingqueryrepresentationsdense; sung-etal-2023-optimizing; 10.1145/3726302.3730160). Minimizing $\mathrm{KL}$ pushes $p_{1}(\cdot\mid z^{(t)})$ to place higher probability where the other distribution does, and to reduce probability where the other distribution is low. However, instead of relying on a heavy cross-encoder, here we utilize feedback from a lightweight bi-encoder whose performance can be on par or even weaker than the primary encoder, and thus the consensus distribution is set as $p_{\mathrm{avg}}$.
Compared to a simple weighted average of scores, GQR operates by updating the query representation of $m_{1}$ rather than defining a new scoring rule. The complementary distribution $p_{\mathrm{avg}}^{(t)}$ is used as a learning signal, and the gradient of $\mathcal{L}^{(t)}$ is computed through $s_{1}(z,d)$, so any change in the ranking must be achievable by moving the query within the primary model’s embedding space. Intuitively, this allows feedback from $m_{2}$ to influence the query in a way that is constrained by $m_{1}$’s notion of similarity, since documents contribute to the update through their embeddings and their relative probabilities. In score-level fusion, each document’s probability is pulled toward the secondary retriever’s distribution by the same fixed amount, so the update is always a uniform weighted average. In GQR, different documents can shift by different magnitudes along a non-linear trajectory dictated by the geometry of that space. In cases where $m_{2}$ is weaker or misaligned, this setup can help incorporate its signal more softly, since it is filtered through $m_{1}$’s representation rather than directly overriding scores.
3 Experiments
3.1 Setup
Task.
Visual document retrieval (mathew2021docvqadatasetvqadocument; mathew2021infographicvqa; li-etal-2024-multimodal-arxiv; zhu2022towards; faysse2025colpali) assumes a corpus of documents, that contain visual elements such as charts, images, and tables, and a set of document-grounded textual queries. The goal is to retrieve the most relevant documents for each query. We conduct experiments on ViDoRe 1 (faysse2025colpali), ViDoRe 2 (mace2025vidore) and the ViDoRe 3 benchmarks, which are established benchmarks for this task. Corpus documents are embedded by encoder models, either directly from page images, or following ingestion of document pages into text (see Appendix D).
Models.
We evaluate a diverse pool of multimodal and textual state-of-the-art retrieval models. The Colpali-based set includes three encoders: Colnomic-Embed-Multimodal-7B (nomicembedmultimodal2025), Jina-Embeddings-v4 (gunther2025jina_v4), and Llama-NemoRetriever-Colembed-3B (xu2025llama_nemo). The set of text models includes Linq-Embed-Mistral (choi2024linq) and Qwen3-Embedding-4B (qwen3embedding), as well as Jina-Embeddings-v4 in its multi-vector textual configuration. This yields $3$ text-based models and $3$ image-based models in total. See Table 4 for details on the models. For the ViDoRe 3 benchmark, we excluded Llama-NemoRetriever-Colembed-3B due to its high computational overhead, which makes evaluation impractical over this larger benchmark.
Metrics and Evaluation.
We use $\mathrm{NDCG}@5$ as the primary metric for our evaluations (we also report $\mathrm{Recall}@5$ in Appendix F.) For each ColPali-based vision-centric model, we test each of the text-centric models as the complementary retriever used for GQR. This yields $9$ GQR pairs in total for ViDoRe 1 and ViDoRe 2 and $6$ for ViDoRe 3, $3$ per ColPali-based model. We also evaluate $4$ different hybrid methods for each vision-text model pair. See Appendix D for technical details about tuning the hyperparameters for ViDoRe 1 and ViDoRe 2.
To simulate real production use cases, we evaluate both tuned and zero-shot settings. We begin with the tuned setting on ViDoRe 1 and ViDoRe 2, and then use the findings from ViDoRe 2 to guide a zero-shot evaluation on ViDoRe 3.
3.2 Comparison to Single-Retriever Pipelines
Table 1: $\mathrm{NDCG}@5$ over ViDoRe 2, by primary and complementary models. Columns show scores by subset and the overall average. Deltas are absolute changes vs. the No refinement row within the same base.
Primary Model GQR complementary model Avg Biomed Lectures Economics ESG Human ESG Full
val $\Delta$ val $\Delta$ val $\Delta$ val $\Delta$ val $\Delta$
Jina (text) $53.4$ $\uparrow$+0.0 $48.6$ $\uparrow$+0.0 $51.4$ $\uparrow$+0.0 $59.5$ $\uparrow$+0.0 $54.1$ $\uparrow$+0.0
Linq-Embed $55.3$ $\uparrow$+0.0 $58.0$ $\uparrow$+0.0 $52.0$ $\uparrow$+0.0 $58.8$ $\uparrow$+0.0 $52.4$ $\uparrow$+0.0
Qwen3 $46.8$ $\uparrow$+0.0 $54.0$ $\uparrow$+0.0 $44.6$ $\uparrow$+0.0 $50.2$ $\uparrow$+0.0 $38.3$ $\uparrow$+0.0
Colnomic-7B
No refinement $60.3$ $\uparrow$+0.0 $64.3$ $\uparrow$+0.0 $54.4$ $\uparrow$+0.0 $68.2$ $\uparrow$+0.0 $54.1$ $\uparrow$+0.0
Jina (text) 63.1 $\uparrow$+2.8 $64.7$ $\uparrow$+0.4 57.0 $\uparrow$+2.6 70.3 $\uparrow$+2.1 $60.2$ $\uparrow$+6.1
Linq-Embed $62.8$ $\uparrow$+2.5 65.4 $\uparrow$+1.1 $56.7$ $\uparrow$+2.3 $67.7$ $\downarrow$1.0 61.2 $\uparrow$+7.1
Qwen3 $61.0$ $\uparrow$+0.7 $61.9$ $\downarrow$2.4 $54.3$ $\downarrow$0.1 $70.2$ $\uparrow$+2.0 $57.5$ $\uparrow$+3.4
Jina (vision)
No refinement $57.2$ $\uparrow$+0.0 $61.6$ $\uparrow$+0.0 $53.5$ $\uparrow$+0.0 $61.7$ $\uparrow$+0.0 $52.0$ $\uparrow$+0.0
Jina (text) $60.7$ $\uparrow$+3.5 $61.7$ $\uparrow$+0.1 $55.3$ $\uparrow$+1.8 $66.9$ $\uparrow$+5.2 58.8 $\uparrow$+6.8
Linq-Embed 61.2 $\uparrow$+4.0 64.7 $\uparrow$+3.1 57.2 $\uparrow$+3.7 $65.7$ $\uparrow$+4.0 $57.1$ $\uparrow$+5.1
Qwen3 $59.8$ $\uparrow$+2.6 $63.2$ $\uparrow$+1.6 $53.6$ $\uparrow$+0.1 67.8 $\uparrow$+6.1 $54.4$ $\uparrow$+2.4
Llama-Nemo
No refinement $63.0$ $\uparrow$+0.0 $63.7$ $\uparrow$+0.0 $56.8$ $\uparrow$+0.0 $74.5$ $\uparrow$+0.0 $56.9$ $\uparrow$+0.0
Jina (text) $64.2$ $\uparrow$+1.2 $64.5$ $\uparrow$+0.8 57.6 $\uparrow$+0.8 $74.2$ $\downarrow$0.3 $60.4$ $\uparrow$+3.5
Linq-Embed 65.2 $\uparrow$+2.2 66.4 $\uparrow$+2.7 $56.8$ $\uparrow$+0.0 74.6 $\uparrow$+0.1 62.8 $\uparrow$+5.9
Qwen3 $63.3$ $\uparrow$+0.3 $65.0$ $\uparrow$+1.3 $55.4$ $\downarrow$1.4 $74.1$ $\downarrow$0.4 $58.7$ $\uparrow$+1.8
Results with Hyperparameter Tuning.
Table 1 reports GQR against the corresponding models on ViDoRe 2. The first block lists the text-only models, which average between $46.8$ for Qwen and $55.3$ for Linq. In each subsequent block, a ColPali-based retriever is fixed, and we show its score alongside the GQR variants, with deltas computed relative to the primary retriever. For Colnomic-7B, the average score rises from $60.3$ to $63.1$ with query refinement from Jina (text) $(+2.8)$ and to $62.8$ with Linq-Embed $(+2.5)$. For Llama-Nemo, the strongest model on ViDoRe 2 to date, GQR improves the average from $63.0$ to $65.2$ with Linq-Embed $(+2.2)$. Notably, the text models clearly underperform the ColPali-based retrievers, yet with GQR the complementary signal they provide boosts performance. This is clearest in the Llama-Nemo versus Qwen3 setting, where GQR delivers a small gain of $\uparrow$+0.3+0.3, despite a $16.2$ point gap in base $\mathrm{NDCG}@5$ (Llama-Nemo $63.0$ vs. Qwen $46.8$). Subsets where GQR harms performance are rare for Colnomic-7B and Llama-Nemo, absent for Jina (vision), and on average across the benchmark GQR consistently improves retrieval quality. We observe similar patterns for the $\mathrm{Recall}@5$ metric (Table 12). On ViDoRe 1 (Table 7 in the Appendix) GQR is generally on par with the base models, yet the benchmark clearly suffers from saturation, with many subset scores reaching $90$ or higher.
Table 2: $\mathrm{NDCG}@5$ over ViDoRe 3, by primary and complementary models. Columns show subset scores, average, and deltas relative to the no-refinement baseline within each primary model.
Primary Model GQR complementary model Avg cs energy fin_en fin_fr hr industrial pharma physics
val $\Delta$ val $\Delta$ val $\Delta$ val $\Delta$ val $\Delta$ val $\Delta$ val $\Delta$ val $\Delta$ val $\Delta$
Jina (text) $51.6$ $\uparrow$+0.0 $66.6$ $\uparrow$+0.0 $60.2$ $\uparrow$+0.0 $54.0$ $\uparrow$+0.0 $40.1$ $\uparrow$+0.0 $54.5$ $\uparrow$+0.0 $42.2$ $\uparrow$+0.0 $55.3$ $\uparrow$+0.0 $40.5$ $\uparrow$+0.0
Linq-Embed $44.2$ $\uparrow$+0.0 $64.6$ $\uparrow$+0.0 $48.5$ $\uparrow$+0.0 $40.1$ $\uparrow$+0.0 $28.7$ $\uparrow$+0.0 $42.6$ $\uparrow$+0.0 $30.1$ $\uparrow$+0.0 $57.7$ $\uparrow$+0.0 $41.4$ $\uparrow$+0.0
Qwen3 $46.0$ $\uparrow$+0.0 $67.6$ $\uparrow$+0.0 $50.1$ $\uparrow$+0.0 $38.9$ $\uparrow$+0.0 $27.3$ $\uparrow$+0.0 $44.6$ $\uparrow$+0.0 $37.8$ $\uparrow$+0.0 $58.1$ $\uparrow$+0.0 $43.9$ $\uparrow$+0.0
Colnomic-7B
No refinement $55.7$ $\uparrow$+0.0 $74.0$ $\uparrow$+0.0 $62.3$ $\uparrow$+0.0 $54.2$ $\uparrow$+0.0 $42.7$ $\uparrow$+0.0 $57.1$ $\uparrow$+0.0 $49.5$ $\uparrow$+0.0 $62.0$ $\uparrow$+0.0 $44.1$ $\uparrow$+0.0
Jina (text) 57.4 $\uparrow$+1.7 $74.8$ $\uparrow$+0.8 64.1 $\uparrow$+1.8 58.6 $\uparrow$+4.4 44.6 $\uparrow$+1.9 59.4 $\uparrow$+2.3 50.8 $\uparrow$+1.3 $62.8$ $\uparrow$+0.8 $44.5$ $\uparrow$+0.4
Linq-Embed $57.1$ $\uparrow$+1.4 $75.1$ $\uparrow$+1.1 $63.4$ $\uparrow$+1.1 $56.9$ $\uparrow$+2.7 $43.9$ $\uparrow$+1.2 $58.1$ $\uparrow$+1.0 50.8 $\uparrow$+1.3 63.4 $\uparrow$+1.4 45.7 $\uparrow$+1.6
Qwen3 $56.7$ $\uparrow$+1.0 75.2 $\uparrow$+1.2 $63.1$ $\uparrow$+0.8 $56.0$ $\uparrow$+1.8 $42.6$ $\downarrow$-0.1 $58.0$ $\uparrow$+0.9 $50.5$ $\uparrow$+1.0 $63.1$ $\uparrow$+1.1 $45.4$ $\uparrow$+1.3
Jina (vision)
No refinement $53.4$ $\uparrow$+0.0 $69.2$ $\uparrow$+0.0 $59.4$ $\uparrow$+0.0 $52.4$ $\uparrow$+0.0 $39.8$ $\uparrow$+0.0 $54.7$ $\uparrow$+0.0 $47.8$ $\uparrow$+0.0 $61.2$ $\uparrow$+0.0 $43.0$ $\uparrow$+0.0
Jina (text) $55.1$ $\uparrow$+1.7 $70.2$ $\uparrow$+1.0 61.5 $\uparrow$+2.1 56.3 $\uparrow$+3.9 42.6 $\uparrow$+2.8 57.1 $\uparrow$+2.4 $48.1$ $\uparrow$+0.3 $61.6$ $\uparrow$+0.4 $43.5$ $\uparrow$+0.5
Linq-Embed 55.4 $\uparrow$+2.0 $72.0$ $\uparrow$+2.8 $61.2$ $\uparrow$+1.8 $54.9$ $\uparrow$+2.5 $41.5$ $\uparrow$+1.7 $56.6$ $\uparrow$+1.9 $48.6$ $\uparrow$+0.8 63.8 $\uparrow$+2.6 $45.1$ $\uparrow$+2.1
Qwen3 $55.0$ $\uparrow$+1.6 72.3 $\uparrow$+3.1 $60.4$ $\uparrow$+1.0 $54.1$ $\uparrow$+1.7 $39.7$ $\downarrow$-0.1 $55.9$ $\uparrow$+1.2 49.3 $\uparrow$+1.5 $63.1$ $\uparrow$+1.9 45.6 $\uparrow$+2.6
Results - Zero Shot.
For ViDoRe 3, we use a single hyperparameter configuration across all subsets and model pairs, a choice informed by our hyperparameter investigation on top of ViDoRe 2 in subsection 4.1. Table 2 shows that GQR improves performance across all model pairs and benchmark subsets, consistently outperforming the no-refinement baseline. For example, on Colnomic-7B it raises the average $\mathrm{NDCG}@5$ from $55.7$ to $57.4$ with Jina as the complementary model (a gain of +1.7+1.7), and delivers per-subset gains as high as +4.4+4.4 on fin_en. Similar improvements appear for Jina (vision), where GQR increases the average from $53.4$ to $55.4$ with Linq-Embed (+2.0+2.0).
3.3 Comparison to hybrid retrieval pipelines
Table 3: Percentage gain in $\mathrm{NDCG}@5$ of hybrid retrieval over the primary retriever on ViDoRe 2. Each cell depicts the average gain over 99 retriever pairs (33 multimodal base retrievers and $3$ text retrievers).
Method Avg Std Per-Subset Gain
Biomed Lectures Economics ESG Human ESG Full
[…]
Table 3 depicts the aggregated performance of the different hybrid retrieval methods on ViDoRe 2, presented as the average percentage gain relative to the base retriever over all pairs (Table 6 in the appendix lists the average absolute values). The Std column captures the standard deviation of performance across the pairs for each method. It shows that the ranking aggregation methods (RRF, Average Ranking) generally lead to a deterioration in performance. All aggregation methods benefit from parameter tuning, with tuned score aggregation methods achieving consistent gains over the base retriever. Table 5 in Appendix F shows the results in zero-shot settings on ViDoRe 3 (see Figure 7 for how the baseline zero-shot configuration is chosen).
Across the benchmarks, GQR outperforms all other hybrid retrieval variants, with the largest average gains over the base multimodal retriever, while also emerging as the most stable positive method with a low standard deviation across pairs.
3.4 Results - Efficiency
Online Querying Latency.
To characterize the test-time costs of GQR and ColPali-based retrievers, we ran latency measurements on a single NVIDIA A100 GPU, measuring document retrieval for randomly sampled queries from each ViDoRe 2 subset (averaging across $100$ runs). Figure 3 depicts the quality-latency trade-offs of the different retrievers, where each point on the graph represents either a base retriever or a retriever with GQR (Appendix F shows full values). The plot illustrates the substantial costs of the strongest base model – Llama-Nemo (orange square, right) – which attains $\mathrm{NDCG}@5=62.9$ at a cost of $2{,}591\,\text{ms}$ per query. Notably, our Colnomic GQR hybrid, with Linq as the refinement model (green diamond, left), reaches $\mathrm{NDCG}@5=62.7$ at $181\,\text{ms}$ ($\approx\!14\times$ faster), and the Colnomic hybrid with Jina (green circle) attains $\mathrm{NDCG}@5=63.0$ at $350\,\text{ms}$ ($\approx\!7\times$ faster), surpassing Nemo. Across base models, applying GQR increases latency by small relative measures with large gains in performance, shifting the Pareto frontier left and upwards. Appendix F shows that GQR is also on the index-storage Pareto frontier.
Figure 3: Latency–quality tradeoff in online querying. The $x$ axis is runtime in milliseconds for a single query, on a log scale, and the $y$ axis is the average evaluation score (NDCG@5). Empty squares indicating the primary retriever alone (without applying GQR).
4 Analysis
4.1 Hyperparameter investigation
GQR relies on two hyperparameters - the learning rate $\alpha$ and the number of optimization steps $T$. We evaluate performance over the Cartesian product $\{1\times 10^{-5},\,5\times 10^{-5},\,10^{-4},\,5\times 10^{-4},\,10^{-3},\,5\times 10^{-3}\}\times\{5,10,15,\ldots,90,95,100\}$ of $\alpha$ and $T$. We run this analysis over 6 model pairs, excluding Llama-NemoRetriever-Colembed-3B due to its computational overhead.
Figure 4: Hyperparameter sweep over GQR’s learning rate $\alpha$ and optimization steps $T$, averaged over six model pairs on ViDoRe 2.
Figure 5: Online latency breakdown of GQR for $T=25$ and $T=50$.
Figure 5 depicts GQR performance on ViDoRe 2 for each combination of $\alpha$ and $T$, averaged over the $6$ model pairs. It shows that high learning rates ($10^{-3},\,5\times 10^{-4}$) are compatible with small $T$ values (1010-$25$), where beyond that performance starts decreasing. Medium learning rates ($10^{-4},\,5\times 10^{-5}$), show consistent improvement and stable optimization, yet require more steps (7575-$100$) to reach higher performance. The smallest ($10^{-5}$) and largest ($5\times 10^{-3}$) learning rates are suboptimal, where the latter even results in performance degradation relative to the primary retriever.
The results capture a tradeoff between latency and stability. Higher learning rates can provide a performance boost faster, but might deteriorate quickly past a certain $T$ value, while medium learning rates are more stable yet require more online optimization steps. Overall, three configurations emerge as the most promising for GQR: (1) $\alpha=10^{-3}$ with $5<T<15$, (2) $\alpha=5\times 10^{-4}$ with $15<T<30$, and (3) $\alpha=10^{-4}$ with $50<T<75$. We leave it to future work to explore methods for optimizing this tradeoff, potentially by incorporating learning-rate scheduling procedures.
4.2 Online latency breakdown
Figure 5 describes the online latency breakdown of GQR, for $T=25$ and $T=50$. The primary model used is Colnomic-Embed-Multimodal-7B and the complementary is Linq-Embed-Mistral. Latencies are collected with an A100 gpu over the ViDoRe 2 benchmark.
The test-time optimization accounts for 50 ms of a 192 ms total latency for $T=25$ and 99 ms of a 241 ms total for $T=50$, averaging roughly 2 ms per optimization step. The procedure is shown here in a sequential form, although with sufficient GPU memory the query encoding and index search for the two retrievers can be parallelized, reducing overall latency.
4.3 Reranker Comparison
Cross-encoder rerankers are widely used in retrieval pipelines, applying full query to document interaction via self-attention to the top-K documents. Similarly to GQR, rerankers operate at test time on a candidate pool returned by a bi-encoder, and thus provide a natural point of comparison. We evaluate GQR against lightonai/MonoQwen2-VL-v0.1, an open-weights multimodal reranker (MonoQwen). Figure 6 illustrates the latency-performance characteristics of GQR against reranking the top-$5$ (on the left) and the top-$10$ (on the right) candidates returned by each retriever. GQR is run with a default $K=10$ configuration and with Linq-Embed as the refinement model. The left plot illustrates that GQR outperforms a top-$5$ reranking pipeline on both the latency and performance axes. Against the top-$10$ reranking pipelines (right), GQR achieves close performance while being $21\times$ faster for Colnomic, $16\times$ faster for Jina (vision), and $2\times$ faster for Nemo. Across both comparisons, GQR remains on the Pareto front and indicates attractive latency performance trade-offs. We report the full results in Table 11.
4.4 GQR Design Choices
Extra search stage.
Prior works on test-time query optimization often run a second index search with the optimized query $z^{(T)}$, aiming to retrieve documents beyond the initial top-$K$ and increase recall (sung-etal-2023-optimizing; 10.1145/3726302.3730160). This extra pass adds latency as it repeats index traversal and candidate generation. We test this modification to GQR where we perform an additional search with the optimized query over the full index, and observe no improvement in performance (Table 16). This suggests that the effects of GQR in our setting are largely confined to the original pool $\mathcal{C}(q)$ of candidate documents.
Choice of objective.
The GQR optimization process minimizes the KL-divergence between the distribution of the primary retriever and a consensus distribution of the two retrievers, $\mathrm{KL}\!(p_{\mathrm{avg}}\|\,p_{1})$. We additionally test two other loss functions: The Jensen-Shannon divergence, $\mathrm{JS}(p_{2}\|\,p_{1})\ =\frac{1}{2}\mathrm{KL}\!(p_{2}\|\,p_{\mathrm{avg}})+\frac{1}{2}\mathrm{KL}\!(p_{1}\|\,p_{\mathrm{avg}})$ and $\mathrm{KL}$ with the target distribution $p_{2}$, i.e., $\mathrm{KL}\!(p_{2}\|\,p_{1})$. We find (Table 18) that these GQR variants generally perform similarly well, indicating that GQR applies across different loss formulations of the two distributions.
We conduct additional experiments on GQR design choices in Appendix E.
5 Related Works
Multimodal Retrieval.
Recent advances in visual document retrieval have been dominated by late-interaction, multi-vector architectures. This paradigm, first introduced to the multimodal domain by ColPali (faysse2025colpali), adapts the ColBERT framework (khattab2020colbert) by treating image patches as visual tokens that interact with textual query tokens via MaxSim operations. Subsequent models built on this foundation – such as Llama-NemoRetriever-ColEmbed (xu2025llama_nemo) and ColNomic-Embed-Multimodal (nomicembedmultimodal2025) – have established state-of-the-art performance by capturing fine-grained interactions that are lost in single-vector representations. However, this performance comes at a significant cost: to achieve their results, these models rely on massively scaled representations, leading to substantial latency and storage overheads that can hinder practical deployment. This pressing trade-off between performance and efficiency motivates our work. Instead of pursuing ever-larger monolithic models, we investigate an alternative direction: enhancing these powerful vision-centric retrievers by fusing their signal with a complementary lightweight text-based encoder at test time.
Hybrid Search.
Numerous works apply hybrid retrieval in the context of combining dense semantic text retrieval with sparse lexical representations (karpukhin-etal-2020-dense; kuzi2020leveraging; luan-etal-2021-sparse; chen2022out). bruch2023analysis conduct a theoretical and empirical analysis of the different ways to perform such dense-sparse fusions. Specifically, they compare RRF (36196) to score-based fusion, and analyze the sensitivity to the choice of tuned weights and normalizations. hsu2025dat propose to set the score-fusion weights dynamically for each query at test time, based on (costly) feedback from an LLM judge. In contrast, GQR departs from these approaches by operating at test time on the representation level of the primary retriever.
Test-time query refinement.
Prior work optimizes query representations during inference in text-only setups using pseudo-relevance feedback, often distilling from a cross-encoder re-ranker to a single-vector dense retriever (yu2021improvingqueryrepresentationsdense; sung-etal-2023-optimizing; 10.1145/3726302.3730160). Cross-encoders provide rich interactions, but incur substantial test-time cost. Here we replace the cross-encoder with a complementary bi-encoder (possibly of a different modality), which preserves low latency while still providing a strong guidance signal.
6 Conclusion
In this work, we introduced Guided Query Refinement (GQR), a novel test-time hybrid retrieval method that refines the query representations of a primary retriever using signals from a complementary one. Unlike traditional hybrid techniques that operate on rankings or scores, GQR leverages representation-level interactions while maintaining efficiency and modularity.
Through extensive experiments on the ViDoRe benchmarks, we demonstrated that GQR consistently improves retrieval performance across diverse model pairs, pushing ColPali-based retrievers to the latency–memory Pareto frontier. Our findings highlight that even weaker retrievers can provide valuable complementary guidance, underscoring the potential for resource-efficient retrieval systems in multimodal large-scale settings.
Appendix A Background
Neural Information Retrieval represents a fundamental shift from traditional lexical matching methods like BM25 (robertson1995okapi_bm25), TF-IDF (salton1988term_tf_idf), and other term-based approaches (zhai2017study_term_based). Unlike these sparse retrieval methods that rely on exact term matches and statistical properties, neural approaches learn dense semantic representations that capture conceptual similarity, enabling retrieval based on meaning rather than just shared vocabulary.
Dense encoders
have transformed information retrieval by learning semantic representations of queries and passages in a shared embedding space. These models typically produce a single dense vector representation for each passage and query, enabling efficient similarity computation through operations like cosine similarity or inner product. Early work by karpukhin-etal-2020-dense introduced Dense Passage Retrieval (DPR), demonstrating that dense representations could outperform traditional sparse methods like BM25 for open-domain question answering. For query, $q$ and passage $p$, The similarity score $S$ is defined as:
$S(q,p)=\frac{q\cdot p}{|q||p|}$ (3)
The BERT-based bi-encoder architecture, where queries and passages are independently encoded using separate or shared BERT models, became the foundational paradigm for neural retrieval. This framework has shaped the field, with numerous dense retrieval models building upon it (reimers2019sentence_bert; xiong2020approximate; qu2020rocketqa; izacard2021unsupervised). The bi-encoder design enables pre-computation of passage embeddings, making real-time retrieval feasible at scale. Recently, advanced models such as Linq-Embed-Mistral (choi2024linq) and Qwen3-Embedding (qwen3embedding) extend this paradigm by leveraging larger language models as the underlying encoder to create more powerful dense representations.
Late-interaction models
compute query-document similarity through more fine-grained interaction mechanisms. Rather than compressing all information into a single vector, these models preserve individual token embeddings for both queries and passages. ColBERT (khattab2020colbert) pioneered this approach with its MaxSim operation, which computes the maximum similarity between each query token and all passage tokens, then aggregates these scores. The MaxSim equation, as defined in ColBERT (see eq. 4)), finds for query token embedding $q_{i}$ the maximum similarity (dot product) with any passage token embedding $p_{j}$. These maximum scores are then summed up to get the final relevance score.
$S(q,p)=\sum_{i=1}^{|Q|}\max_{j=1}^{|P|}q_{i}\cdot p_{j}^{T}$ (4)
This approach balances effectiveness and efficiency, as passage representations can still be pre-computed and indexed while the token-level matching captures finer details than single-vector approaches. Subsequent work improved efficiency and retrieval quality through various optimizations (santhanam2021colbertv2; santhanam2022plaid).
Contrastive training
forms the backbone of modern retrieval model optimization. These methods learn representations by pulling positive query-passage pairs closer while pushing negative pairs apart in the embedding space. The InfoNCE loss (oord2018representation_infonce), which maximizes the similarity of positive pairs relative to negative ones through a softmax-like normalization, has become the dominant training objective across retrieval architectures.
$\mathcal{L}=-\log\frac{\exp(\text{sim}(q,p^{+})/\tau)}{\exp(\text{sim}(q,p^{+})/\tau)+\sum_{i=1}^{k}\exp(\text{sim}(q,p_{i}^{-})/\tau)}$ (5)
Cross-encoder Re-rankers
represent a different paradigm where query and passage are jointly encoded, enabling deep self-attention between their representations. Unlike bi-encoders that independently encode queries and passages, cross-encoders process query-passage pairs through a single model, allowing full attention across all tokens. This joint encoding is computationally expensive, making it impractical for first-stage retrieval over large corpora. However, cross-encoders excel as re-rankers, refining the top-K results from efficient first-stage retrievers. Examples include MonoBERT (nogueira2019multi_mono_bert) for text retrieval and MonoQwen; wasserman2025docrerank for visual document retrieval.
A.1 Visual Document Retrieval
Verbalization-based methods
were the dominant approach before the advent of end-to-end vision models. These pipelines convert visual documents into text through various techniques: traditional Optical Character Recognition (OCR) tools like docling (auer2024docling) extract printed text, while Vision-Language Models (VLMs) can generate textual descriptions of visual elements such as charts, diagrams, and infographics. After verbalization, these methods apply standard text retrieval techniques to the extracted content. While verbalization-based approaches can leverage powerful text-only retrieval models, they inherently lose spatial relationships and visual context during the text extraction process.
ColPali architectures
introduced a new approach to visual document retrieval by directly encoding document images without intermediate text extraction. These VLM-based embedding models transform pre-trained generative Vision-Language Models (such as PaliGemma, beyer2024paligemma and Qwen-VL, wang2024qwen2) into multi-vector embedding models optimized for retrieval.
ColPali (faysse2025colpali) was the first model to introduce this approach, building upon PaliGemma (beyer2024paligemma) and adapting the late-interaction framework to vision-language models by treating image patches as visual tokens that interact with textual query tokens through MaxSim operations. ColPali provides native text-query support due to its VLM-based design. Queries remain text, are encoded by the model’s language tower, and are matched directly against visual page/passage tokens, eliminating any OCR at query time. Training is contrastive, typically InfoNCE-style with in-batch/hard negatives to align query tokens with relevant visual tokens. This architecture preserves spatial layout information and visual features that verbalization-based methods discard. Following ColPali’s success, subsequent models have adopted and extended this ColPali paradigm by leveraging different base VLMs: Llama-NemoRetriever-ColEmbed (xu2025llama_nemo), Colnomic-Embed-Multimodal (nomicembedmultimodal2025), Granite-Vision-Embedding (granite_vision_embedding) and Jina-Embeddings-v4 (gunther2025jina_v4).
Appendix B Model Information
Table 4 depicts the details of the models evaluated in our work. Storage assumes a 16-bit representation and follows official reports (xu2025llama_nemo). Jina-Embeddings-v4 supports both text and image document representations, single-vector and multi-vector retrieval, and flexible embedding sizes via a Matryoshka scheme (kusupati2022matryoshka). In its multi-vector textual configuration, the number of vectors per page for Jina varies between pages.
Table 4: Different retrieval models evaluated in our work. The modality column describes the representation used for the documents – page-level text, or page-level images.
Model Size # Vectors per Page Token Dim # Floats per Page Storage per 1M Docs (GB) Page Modality
Linq-Embed-Mistral 7.1B 1 4096 4096 7.63 Text
Qwen3-Embedding-4B 4B 1 2560 2560 4.77 Text
Jina-Embeddings-v4 (Text) 3.8B 128 Text
Jina-Embeddings-v4 (Image) 3.8B 767 128 98176 182.87 Image
Colnomic-Embed-Multimodal-7B 7B 767 128 98176 182.87 Image
Llama-NemoRetriever-Colembed-3B 4.4B 1802 3072 5535744 10311.13 Image
Appendix C Additional Definitions
Average Ranking
Computes the average rank across retrievers.
$\mathrm{AvgRank}(d)=\frac{1}{M}\sum_{m=1}^{M}\mathrm{rank}_{m}(d)$ (6)
Min-Max normalization
$\tilde{s}_{m}(q,d_{i})=\frac{s_{m}(q,d_{i})-\min_{d_{j}\in\pi_{m}(q)}s_{m}(q,d_{j})}{\max_{d_{j}\in\pi_{m}(q)}s_{m}(q,d_{j})-\min_{d_{j}\in\pi_{m}(q)}s_{m}(q,d_{j})+\varepsilon},$ (7)
with a small $\varepsilon$ for numerical stability.
Softmax normalization
$\tilde{s}_{m}(q,d_{i})=\frac{\exp\!\left(s_{m}(q,d_{i})\right)}{\sum_{d_{j}\in\pi_{m}(q)}\exp\!\left(s_{m}(q,d_{j})\right)}.$ (8)
If $d_{i}\notin\pi_{m}(q)$ we set $\tilde{s}_{m}(q,d_{i})=0$
Average Ranking - Tuned (for M=2)
$\mathrm{AvgRank}(d)=\alpha\,\ {rank}_{m_{1}}(d)+(1-\alpha)\,\ {rank}_{m_{2}}(d)$ (9)
RRF - Tuned (for M=2)
$\mathrm{RRF}(d)=\frac{2\alpha}{\,\kappa+\mathrm{rank}_{m_{1}}(d)\,}+\frac{2(1-\alpha)}{\,\kappa+\mathrm{rank}_{m_{2}}(d)\,}$ (10)
Score aggregation - Tuned (for M=2)
$\mathrm{Score}(q,d_{i})=\alpha\,\ \tilde{s}_{m_{1}}(q,d_{i})+(1-\alpha)\,\ \tilde{s}_{m_{2}}(q,d_{i})$ (11)
KL divergence.
$\mathrm{KL}\!\big(P\|Q\big)=\sum_{d\in\mathcal{C}(q)}P(d)\,\log\!\frac{P(d)}{Q(d)}\,$, (12)
where $P$ and $Q$ are distributions over $\mathcal{C}(q)$.
Appendix D Implementation Details
Offline indexing.
For each model and dataset, we construct an offline index of page-level document representations. For ColPali-based models, multi-page documents are rendered as page images and encoded directly. For text-only embedding models, for the ViDoRe 1 and ViDoRe 2 datasets an ingestion pipeline converts each page to text. We use Docling [^6] https://docling-project.github.io/docling/(Docling) to ingest the images. Docling is an open library providing OCR capabilities combined with document layout analysis, allowing us to recover page content via simple function calls. The resulting text is stored alongside the page images without any chunking, ensuring consistent alignment between visual and textual page representations across the datasets. We run the document converter from Docling v2.342.34 using default parameters. For ViDoRe 3, we use the official markdown of the benchmark as the text input for the embedding models.
For Jina-Embeddings-v4 (gunther2025jina_v4), a single model accepts both image and text input documents. We thus use it in both a vision configuration and in a text configuration, denoted Jina (Vision) and Jina (Text), both running in a multi-vector setting. Jina (Text) is thus used to test the applicability of GQR where a multi-vector architecture is used as the the GQR complementary encoder.
Hyperparameters
For RRF we set $\kappa=60$, a common default (chen2022out; 36196). For tuning the weighted hybrid baselines and GQR, we follow previous works (sung-etal-2023-optimizing; 10.1145/3726302.3730160; bruch2023analysis) and rely on an in-domain development set of queries. We reserve 10% of each subset and tune the hyperparameters, $T$, and $\alpha$ for each, selecting by development set $\mathrm{NDCG}@5$ performance. The splits are fixed for all experiments. For the weighted hybrid variants, the weight $\alpha$ is tuned over $\{0.1,0.2,\ldots,0.9\}$ for each subset. For GQR, the learning rates are $\{1\times 10^{-5},\,5\times 10^{-5},\,10^{-4},\,5\times 10^{-4},\,10^{-3},\,5\times 10^{-3}\}$, and we consider $T\in\{10,25,50\}$. We report the selected hyperparameters for all methods in Tables 20-25. We opted for Adam (kingma2015adam) as the optimizer for GQR, based on preliminary experiments comparing different optimizer choices. We use $K=10$ across all retrievers and methods.
Figure 6: Latency–quality tradeoff in online querying. The $x$ axis is runtime in milliseconds for a single query, on a log scale, and the $y$ axis is the average evaluation score (NDCG@5). Marker color encodes the primary model; marker shape encodes the usage of GQR or reranking with top 5 (left) or top 10 (right) candidates. The dashed lines represent the Pareto frontier.
Algorithm 1 Guided Query Refinement (GQR)
1:Query $q$, primary encoder $m_{1}$, complementary encoder $m_{2}$, iterations $T$, step size $\alpha$, top-$K$ value $K$
[…]
6:for $t=0$ to T1T-1 do
7:  $p_{1}(d_{i}\mid z^{(t)})\leftarrow\operatorname{softmax}\!\big(s_{1}(z^{(t)},d_{i})\big)$ for $d_{i}\in\mathcal{C}(q)$ $\triangleright$ Primary distribution on $\mathcal{C}(q)$
8:  $p_{\mathrm{avg}}(d_{i}\mid z^{(t)})\leftarrow\tfrac{1}{2}\big(p_{1}(d_{i}\mid z^{(t)})+p_{2}(d_{i}\mid e_{2}^{q})\big)$ $\triangleright$ Consensus (average) distribution
9:  $\mathcal{L}_{\mathrm{KL}}\leftarrow\operatorname{KL}\!\big(p_{\mathrm{avg}}(d_{i}\mid z^{(t)})\,\|\,p_{1}(d_{i}\mid z^{(t)})\big)$ $\triangleright$ Compute the loss
10:  $z^{(t+1)}\leftarrow z^{(t)}-\alpha\,\nabla_{z^{(t)}}\mathcal{L}_{\mathrm{KL}}$ $\triangleright$ Gradient step on the query representation
11:end for
[…]
14:return $\mathcal{R}(q)$
Appendix E Additional Design Choices for GQR
Candidate pool policy.
In our implementation the pool $\mathcal{C}(q)$ of candidate documents used for query refinement is a union of the top-K documents from the primary and complementary retrievers. One alternative is to opt for a reranker-like setup, where the only documents considered are those initially returned in the top-K of the primary retriever. Our analysis shows (Table 17) that this modification does not have a consistent positive or negative effect compared to a union pool of candidates. While the complementary dense retriever’s index search is relatively quick, this configuration can be adopted in sensitive deployments of GQR for additional latency gains.
Primary and complementary roles
We evaluate text-centric and vision-centric retrievers as both primary and complementary retrievers within GQR, reporting results in Table 19. Across model pairs, both role assignments improve over the base retrievers. The alternative role configuration yields larger gains relative to the primary encoder alone, while the default GQR attains the strongest absolute score on ViDoRe 2.
Appendix F Additional Results
Figure 7: Baseline results on ViDoRe 2 across different values of the weight $\alpha$, averaged over 66 model pairs. It shows that the optimal weight for each method varies. Generally the softmax variant emerges as the most stable, while the min-max variant reaches the best performance if tuned correctly. The values for the zero-shot configuration for each baseline in Table 5 were selected based on this plot.
Table 5: Percentage gain in $\mathrm{NDCG}@5$ of hybrid retrieval over the primary retriever on ViDoRe 3. Each cell reports the average gain and standard deviation over 66 retriever pairs (22 multimodal base retrievers and $3$ text retrievers). The Hyperparameters column indicates which hyperparameters were selected for zero-shot evaluation.
Method Avg Std Hyperparameters
[…]
Table 6: $\mathrm{NDCG}@5$ for hybrid retrieval on ViDoRe 2. Results are averaged across 99 model pairs.
Method Avg Biomed Lectures Economics ESG Human ESG Full
Average Ranking 58.1 59.0 55.1 63.6 54.5
RRF 58.2 59.1 55.0 63.8 55.0
Score Aggregation (Min-Max) 60.2 61.3 55.8 66.7 56.8
Score Aggregation (Softmax) 60.9 62.6 55.3 68.4 57.2
Average Ranking - Tuned 59.8 61.3 55.6 66.2 56.0
RRF - Tuned 59.9 61.4 55.5 65.3 57.3
Score Aggregation (Min-Max) - Tuned 62.1 63.9 56.5 70.0 58.0
Score Aggregation (SoftMax) - Tuned 61.6 63.4 55.4 69.5 58.2
Guided Query Refinement (GQR) 62.3 64.2 56.0 70.2 59.0
Table 7 presents the results on ViDoRe 1. It is noticeable that this benchmark suffers from saturation, with many subset scores reaching $90$ or higher (and indeed this was the direct motivation for the release of ViDoRe 2, mace2025vidore). Nevertheless, our method does not harm performance, in contrast to other hybrid retrieval methods, as seen in Tables 8 and 9.
Storage.
Figure 9 in the Appendix shows that the added storage in GQR is modest. As in the latency plot, GQR dominates the strongest base model by a wide margin. The Llama-Nemo index represents each document with 10.6 MB of memory, whereas the Colnomic hybrid, using Linq as the refinement model, nearly matches its quality with $0.20$ MB per document $(\approx 54\times$ less). Using Jina, GQR surpasses Nemo while requiring $0.37$ MB per document $(\approx 28\times$ less).
Appendix G Analyzing Per-query Dynamics
As we explain in subsection 2.4, the dynamics in GQR differ from a simple weighted average of scores – the distribution changes are constrained by the geometry of the primary model’s embedding space, and different documents can shift by different magnitudes. To demonstrate this more directly, in Figure 8 we visualize the changes in document ranks for a given query as a function of the number of GQR steps $t$, and as function of the weight $\alpha$ of the Score Aggregation (Min-Max) baseline. As can be seen in the figure, for some queries the dynamics of GQR and score aggregation are relatively similar, whereas for others the effects and dynamics are starkly different. Notably, while score aggregation operates across all candidate documents, the effects of GQR are often focused on a specific subset of documents.
Figure 8: Query-level dynamics of GQR versus score aggregation. The heat maps depict examples of how the document ranks change as a function of either the GQR optimization step $t$ (center left panels), or the weight $\alpha$ of a linear score aggregation baseline (Score Aggregation - Min-Max, center right panels). For both methods we see a gradual progression from the ranks of the primary retriever $m_{1}$ (left) towards those of the complementary retriever $m_{2}$ (right). However, the dynamics differ across queries, as well as across documents within a query. Examples are taken from the ESG subset of ViDoRe 2, for the model pair Colnomic-Embed-Multimodal-7B,Linq-Embed-Mistral (using a GQR learning rate of $10^{-4}$).
Table 7: $\mathrm{NDCG}@5$ over ViDoRe 1, by primary and complementary models. Deltas are absolute changes vs. the No refinement row within the same base.
Primary Model GQR complementary model Avg ArXivQA DocVQA InfoVQA TabFQuAD TAT DQA ShiftProj SynDocQA AI SynDocQA Energy SynDocQA Gov SynDocQA Health
val $\Delta$ val $\Delta$ val $\Delta$ val $\Delta$ val $\Delta$ val $\Delta$ val $\Delta$ val $\Delta$ val $\Delta$ val $\Delta$ val $\Delta$
jina (text) $80.9$ $\uparrow$+0.0 $42.9$ $\uparrow$+0.0 $43.7$ $\uparrow$+0.0 $69.2$ $\uparrow$+0.0 $97.6$ $\uparrow$+0.0 $82.4$ $\uparrow$+0.0 $85.6$ $\uparrow$+0.0 $100.0$ $\uparrow$+0.0 $94.4$ $\uparrow$+0.0 $98.9$ $\uparrow$+0.0 $94.4$ $\uparrow$+0.0
Linq-Embed $73.8$ $\uparrow$+0.0 $45.3$ $\uparrow$+0.0 $36.0$ $\uparrow$+0.0 $78.2$ $\uparrow$+0.0 $90.5$ $\uparrow$+0.0 $45.9$ $\uparrow$+0.0 $75.6$ $\uparrow$+0.0 $94.4$ $\uparrow$+0.0 $88.9$ $\uparrow$+0.0 $90.0$ $\uparrow$+0.0 $93.3$ $\uparrow$+0.0
Qwen3 $76.4$ $\uparrow$+0.0 $48.2$ $\uparrow$+0.0 $36.9$ $\uparrow$+0.0 $77.2$ $\uparrow$+0.0 $97.2$ $\uparrow$+0.0 $48.7$ $\uparrow$+0.0 $77.8$ $\uparrow$+0.0 $97.8$ $\uparrow$+0.0 $91.1$ $\uparrow$+0.0 $97.8$ $\uparrow$+0.0 $91.1$ $\uparrow$+0.0
Colnomic-7B
No refinement 89.8 $\uparrow$+0.0 88.6 $\uparrow$+0.0 $60.2$ $\uparrow$+0.0 92.4 $\uparrow$+0.0 $96.7$ $\uparrow$+0.0 $81.2$ $\uparrow$+0.0 88.7 $\uparrow$+0.0 99.6 $\uparrow$+0.0 95.9 $\uparrow$+0.0 95.1 $\uparrow$+0.0 99.2 $\uparrow$+0.0
Jina (text) $89.7$ $\downarrow$-0.1 $86.9$ $\downarrow$-1.7 61.3 $\uparrow$+1.1 92.4 $\uparrow$+0.0 $96.7$ $\uparrow$+0.0 $81.2$ $\uparrow$+0.0 88.7 $\uparrow$+0.0 99.6 $\uparrow$+0.0 95.9 $\uparrow$+0.0 95.1 $\uparrow$+0.0 99.2 $\uparrow$+0.0
Linq-Embed 89.8 $\uparrow$+0.0 88.6 $\uparrow$+0.0 $60.7$ $\uparrow$+0.5 $92.0$ $\downarrow$-0.4 $96.7$ $\uparrow$+0.0 81.3 $\uparrow$+0.1 88.7 $\uparrow$+0.0 99.6 $\uparrow$+0.0 95.9 $\uparrow$+0.0 95.1 $\uparrow$+0.0 99.2 $\uparrow$+0.0
Qwen3 89.8 $\uparrow$+0.0 $88.4$ $\downarrow$-0.2 $59.9$ $\downarrow$-0.3 92.4 $\uparrow$+0.0 97.2 $\uparrow$+0.5 $81.2$ $\uparrow$+0.0 88.7 $\uparrow$+0.0 99.6 $\uparrow$+0.0 95.9 $\uparrow$+0.0 95.1 $\uparrow$+0.0 99.2 $\uparrow$+0.0
Jina (vision)
No refinement 89.9 $\uparrow$+0.0 $88.6$ $\uparrow$+0.0 62.4 $\uparrow$+0.0 $92.0$ $\uparrow$+0.0 $96.2$ $\uparrow$+0.0 $78.4$ $\uparrow$+0.0 91.5 $\uparrow$+0.0 99.2 $\uparrow$+0.0 96.1 $\uparrow$+0.0 96.5 $\uparrow$+0.0 98.5 $\uparrow$+0.0
Jina (text) $89.8$ $\downarrow$-0.1 $87.9$ $\downarrow$-0.7 62.4 $\uparrow$+0.0 $92.0$ $\uparrow$+0.0 $96.2$ $\uparrow$+0.0 78.5 $\uparrow$+0.1 91.5 $\uparrow$+0.0 99.2 $\uparrow$+0.0 96.1 $\uparrow$+0.0 96.5 $\uparrow$+0.0 $97.7$ $\downarrow$-0.8
Linq-Embed $89.8$ $\downarrow$-0.1 $88.8$ $\uparrow$+0.2 $61.0$ $\downarrow$-1.4 92.1 $\uparrow$+0.1 $96.2$ $\uparrow$+0.0 $78.4$ $\uparrow$+0.0 91.5 $\uparrow$+0.0 99.2 $\uparrow$+0.0 96.1 $\uparrow$+0.0 96.5 $\uparrow$+0.0 98.5 $\uparrow$+0.0
Qwen3 $89.6$ $\downarrow$-0.3 88.9 $\uparrow$+0.3 $59.8$ $\downarrow$-2.6 $92.0$ $\uparrow$+0.0 $96.2$ $\uparrow$+0.0 $78.4$ $\uparrow$+0.0 91.5 $\uparrow$+0.0 99.2 $\uparrow$+0.0 96.1 $\uparrow$+0.0 96.5 $\uparrow$+0.0 $97.7$ $\downarrow$-0.8
Llama-Nemo
No refinement 91.0 $\uparrow$+0.0 88.0 $\uparrow$+0.0 66.2 $\uparrow$+0.0 94.9 $\uparrow$+0.0 $96.7$ $\uparrow$+0.0 $81.0$ $\uparrow$+0.0 89.9 $\uparrow$+0.0 100.0 $\uparrow$+0.0 $96.3$ $\uparrow$+0.0 $97.7$ $\uparrow$+0.0 99.2 $\uparrow$+0.0
Jina (text) 91.0 $\uparrow$+0.0 88.0 $\uparrow$+0.0 $65.9$ $\downarrow$-0.3 $94.7$ $\downarrow$-0.2 97.0 $\uparrow$+0.3 81.5 $\uparrow$+0.5 $89.6$ $\downarrow$-0.3 100.0 $\uparrow$+0.0 $96.3$ $\uparrow$+0.0 $97.7$ $\uparrow$+0.0 99.2 $\uparrow$+0.0
Linq-Embed 91.0 $\uparrow$+0.0 88.0 $\uparrow$+0.0 $65.8$ $\downarrow$-0.4 $94.5$ $\downarrow$-0.4 $96.9$ $\uparrow$+0.2 $80.9$ $\downarrow$-0.1 $89.7$ $\downarrow$-0.2 100.0 $\uparrow$+0.0 96.7 $\uparrow$+0.4 98.1 $\uparrow$+0.4 99.2 $\uparrow$+0.0
Qwen3 $90.8$ $\downarrow$-0.2 $87.2$ $\downarrow$-0.8 $65.7$ $\downarrow$-0.5 94.9 $\uparrow$+0.0 $96.6$ $\downarrow$-0.1 $80.8$ $\downarrow$-0.2 $89.7$ $\downarrow$-0.2 100.0 $\uparrow$+0.0 $95.8$ $\downarrow$-0.5 98.1 $\uparrow$+0.4 99.2 $\uparrow$+0.0
Table 8: $\mathrm{NDCG}@5$ for hybrid retrieval on ViDoRe 1. Results are averaged across 99 retriever pairs.
Method Avg arXivQA DocVQA InfoVQA TabFQuAD TATDQA ShiftProj SynthAI SynthEnergy SynthGov SynthHealth
Average Ranking 77.8 55.6 44.0 76.2 93.4 61.1 78.5 96.0 89.8 91.0 92.7
RRF 78.0 53.8 44.1 76.5 93.5 61.7 79.0 96.1 90.3 91.4 93.3
Score Aggregation (Min-Max) 84.4 74.4 53.0 85.5 95.4 69.4 83.2 97.9 94.0 95.4 95.9
Score Aggregation (Softmax) 88.6 83.3 57.7 91.0 96.5 78.3 88.2 99.6 96.0 96.6 98.6
Average Ranking - Tuned 85.5 79.5 57.3 87.0 95.5 74.4 86.0 97.8 88.3 92.3 96.9
RRF - Tuned 84.5 76.6 55.8 85.9 95.7 73.4 84.4 97.6 88.0 92.3 95.2
Score Aggregation (Min-Max) - Tuned 88.5 88.0 62.2 92.6 96.1 79.6 88.4 97.7 88.9 93.9 98.0
Score Aggregation (SoftMax) - Tuned 89.4 87.7 62.4 92.4 96.5 80.2 86.9 99.1 95.3 96.0 97.6
Guided Query Refinement (GQR) 90.1 88.1 62.5 93.0 96.6 80.2 90.0 99.6 96.1 96.5 98.8
Table 9: Percentage gain, in $\mathrm{NDCG}@5$, of hybrid retrieval over the primary retriever for ViDoRe 1. Each cell depicts average gain over 99 retriever pairs (33 multimodal base retrievers $\times$ 33 text retrievers).
Method Avg arXivQA DocVQA InfoVQA TabFQuAD TATDQA ShiftProj SynthAI SynthEnergy SynthGov SynthHealth
[…]
Table 10: Performance, latency (ms per query), and memory (MB per document) by primary and complementary models.
Primary Model Complementary Model Performance Latency Latency Diff Memory (MB)
Colnomic-7b No refinement 60.25 115.98 0.20
Colnomic-7b Linq-Embed 62.75 181.21 65.23 0.20
Colnomic-7b Qwen3 60.98 196.16 80.15 0.20
Colnomic-7b Jina (text) 63.05 350.13 194.15 0.39
Jina (vision) No refinement 57.20 153.45 0.20
Jina (vision) Linq-Embed 61.18 213.97 60.5 0.20
Jina (vision) Qwen3 59.75 233.06 79.61 0.20
Jina (vision) Jina (text) 60.68 394.64 241.19 0.39
Llama-Nemo No refinement 62.98 2591.14 11.07
Llama-Nemo Linq-Embed 65.15 2674.84 83.7 11.08
Llama-Nemo Qwen3 63.30 2712.12 120.98 11.08
Llama-Nemo Jina (text) 64.18 2934.61 343.47 11.27
Figure 9: Storage–quality tradeoff. The $x$ axis is memory in MB, on a log scale, and the $y$ axis is the average evaluation score (NDCG@5). Marker color encodes the primary retriever; marker shape encodes the GQR complementary retriever, with squares indicating the primary retriever alone (without applying GQR).
Table 11: Performance and end-to-end latency of reranking pipelines against GQR. Rows are grouped by reranker candidate size $k$. A dedicated block reports GQR.
Reranking $k$ Retriever Latency NDCG@5 $\mathrm{Recall}@5$
No-reranking Colnomic-7B (multi) 115.98 60.25 57.32
Jina (vision, multi) 153.45 57.20 56.17
Llama Nemo 3B (multi) 2591.14 62.97 59.75
GQR Colnomic-7B (multi) 181.21 62.75 58.0
Jina (vision, multi) 213.97 61.0 57.6
Llama Nemo 3B (multi) 2674.83 65.15 60.1
5 Colnomic-7B (multi) 1823.03 62.12 57.32
Jina (vision, multi) 1860.714 60.70 56.17
Llama Nemo 3B (multi) 4332.55 64.72 59.75
10 Colnomic-7B (multi) 3586.809 64.37 59.92
Jina (vision, multi) 3585.946 63.27 58.8
Llama Nemo 3B (multi) 6027.078 66.10 61.8
20 Colnomic-7B (multi) 7035.953 65.07 60.27
Jina (vision, multi) 7251.81 64.10 59.6
Llama Nemo 3B (multi) 9470.134 65.77 61.02
Table 12: $\mathrm{Recall}@5$ on ViDoRe 2, by primary and complementary models.
Avg Biomed Lectures Economics ESG Human ESG Full
Primary Model Complementary Model
Jina-Embeddings (Text) 50.2 50.7 26.3 68.1 55.5
Linq-Embed 50.1 60.3 25.7 62.7 51.8
Qwen3-Embedding 44.6 57.5 25.5 53.6 41.9
Colnomic-Embed 57.3 66.9 30.9 74.2 57.3
Jina-Embeddings (Text) 58.7 67.1 30.1 74.9 62.7
Linq-Embed 58.0 68.3 29.7 72.6 61.5
Qwen3-Embedding 58.7 65.1 30.8 77.3 61.4
Jina-Embeddings 56.2 64.2 29.6 71.8 59.1
Jina-Embeddings (Text) 57.5 64.2 27.7 75.9 62.1
Linq-Embed 57.6 66.9 29.4 72.3 61.7
Qwen3-Embedding 56.9 66.2 29.4 71.7 60.2
Llama-Nemoretriever 59.8 66.5 30.7 80.1 61.7
Jina-Embeddings (Text) 60.2 65.9 30.5 80.1 64.1
Linq-Embed 60.1 68.4 28.4 79.5 64.1
Qwen3-Embedding 59.9 67.1 30.4 79.3 62.7
Table 13: $\mathrm{Recall}@5$ for hybrid retrieval on ViDoRe 2. Results are averaged across 99 retriever pairs.
Method Avg Biomed Lectures Economics ESG Human ESG Full
Average Ranking 54.3 61.4 29.1 69.4 57.3
RRF 54.6 61.5 29.1 69.7 58.0
Score Aggregation (Min-Max) 56.7 64.7 29.7 72.3 59.9
Score Aggregation (Softmax) 57.3 65.3 29.9 73.7 60.4
Average Ranking - Tuned 56.4 64.9 29.5 71.5 59.7
RRF - Tuned 56.3 64.6 29.5 70.5 60.7
Score Aggregation (Min-Max) - Tuned 58.5 66.9 30.3 75.5 61.4
Score Aggregation (SoftMax) - Tuned 57.9 65.8 29.6 74.5 61.6
Guided Query Refinement (GQR) 58.6 66.6 29.6 76.0 62.3
Table 14: $\mathrm{Recall}@5$ on ViDoRe 1, by primary and complementary models.
Avg arXivQA DocVQA InfoVQA TabFQuAD TATDQA ShiftProj SynthAI SynthEnergy SynthGov SynthHealth
Model Complementary Model
Jina-Embeddings (Text) 42.9 80.9 43.7 69.2 85.6 100.0 94.4 98.9 94.4 97.6 82.4
Linq-Embed 45.3 73.8 36.0 78.2 75.6 94.4 88.9 90.0 93.3 90.5 45.9
Qwen3-Embedding 48.2 76.4 36.9 77.2 77.8 97.8 91.1 97.8 91.1 97.2 48.7
Colnomic-Embed 93.1 93.7 66.2 95.5 95.6 100.0 97.8 100.0 100.0 98.8 89.5
Jina-Embeddings (Text) 90.0 93.4 67.2 95.5 95.6 100.0 97.8 100.0 100.0 98.8 89.4
Linq-Embed 93.1 93.7 66.5 95.3 95.6 100.0 97.8 100.0 100.0 98.8 89.7
Qwen3-Embedding 93.1 93.8 66.3 96.0 95.6 100.0 97.8 100.0 100.0 99.2 89.6
Jina-Embeddings 92.2 94.2 71.0 95.4 97.8 100.0 97.8 100.0 100.0 98.8 88.5
Jina-Embeddings (Text) 90.9 94.0 71.0 95.4 97.8 100.0 97.8 100.0 100.0 98.8 88.5
Linq-Embed 92.0 94.0 69.6 95.5 97.8 100.0 97.8 100.0 100.0 98.8 88.7
Qwen3-Embedding 92.2 93.9 68.7 95.4 97.8 100.0 97.8 100.0 100.0 98.8 88.5
Llama-Nemoretriever 92.4 94.7 73.0 98.0 98.9 100.0 96.7 100.0 100.0 99.6 88.8
Jina-Embeddings (Text) 92.4 94.6 72.7 98.0 97.8 100.0 96.7 100.0 100.0 99.6 89.0
Linq-Embed 92.4 94.8 72.4 97.6 98.9 100.0 97.8 100.0 100.0 99.6 88.9
Qwen3-Embedding 91.6 94.5 71.5 98.2 98.9 100.0 96.7 100.0 100.0 98.8 88.8
Table 15: $\mathrm{Recall}@5$ for hybrid retrieval on ViDoRe 1. Results are averaged across 99 retriever pairs.
Method Avg arXivQA DocVQA InfoVQA TabFQuAD TATDQA ShiftProj SynthAI SynthEnergy SynthGov SynthHealth
Average Ranking 84.1 63.9 51.8 81.5 97.1 71.6 87.8 99.3 93.7 98.0 96.7
RRF 83.8 59.5 51.3 81.8 97.2 72.4 88.4 99.0 93.7 98.3 96.7
Score Aggregation (Min-Max) 91.5 88.4 63.5 93.1 98.6 83.0 92.3 100.0 97.6 99.9 98.4
Score Aggregation (Softmax) 92.5 87.4 63.5 94.5 99.1 87.0 96.1 100.0 97.3 100.0 100.0
Average Ranking - Tuned 92.9 92.2 68.7 95.6 98.8 87.6 95.2 99.9 93.3 98.2 100.0
RRF - Tuned 91.5 90.9 66.2 93.5 98.8 85.0 92.4 99.5 92.8 98.2 97.2
Score Aggregation (Min-Max) - Tuned 93.5 92.3 69.3 96.3 99.0 88.7 96.6 99.5 93.8 99.1 100.0
Score Aggregation (SoftMax) - Tuned 93.5 91.4 69.4 95.8 99.0 88.9 94.7 99.9 97.1 99.6 98.8
Guided Query Refinement (GQR) 94.1 92.0 69.5 96.3 99.0 89.0 97.3 100.0 97.6 100.0 100.0
Table 16: Effect of extra index search on GQR $\mathrm{NDCG}@5$ performance, over ViDoRe 2.
Avg Biomed Lectures Economics ESG Human ESG Full
Model Complementary Model Variant
Colnomic-Embed Jina-Embeddings GQR 63.0 64.7 57.0 70.3 60.2
GQR + Search 63.1 64.7 57.1 70.3 60.2
Linq-Embed GQR 62.8 65.4 56.7 67.7 61.2
GQR + Search 62.7 65.4 56.5 67.9 61.0
Qwen3-Embedding GQR 61.0 61.9 54.3 70.2 57.5
GQR + Search 61.0 61.7 54.3 70.6 57.5
Jina-Embeddings Jina-Embeddings GQR 60.7 61.7 55.3 66.9 58.8
GQR + Search 60.7 61.7 55.2 66.9 58.9
Linq-Embed GQR 61.2 64.7 57.2 65.7 57.1
GQR + Search 61.0 64.7 57.2 65.0 57.2
Qwen3-Embedding GQR 59.8 63.2 53.6 67.8 54.4
GQR + Search 59.8 63.2 53.6 67.8 54.4
Llama-Nemoretriever Jina-Embeddings GQR 64.2 64.5 57.6 74.2 60.4
GQR + Search 64.1 64.4 57.6 74.2 60.4
Linq-Embed GQR 65.1 66.4 56.8 74.6 62.8
GQR + Search 65.3 66.5 57.2 74.6 62.8
Qwen3-Embedding GQR 63.3 65.0 55.4 74.1 58.7
GQR + Search 63.3 64.8 55.4 74.1 58.7
Table 17: Effect of candidate pool on GQR $\mathrm{NDCG}@5$ performance, over ViDoRe 2.
Avg Biomed Lectures Economics ESG Human ESG Full
Model Complementary Model Variant
Colnomic-Embed Jina-Embeddings GQR 63.0 64.7 57.0 70.3 60.2
GQR (Top-$K$ only) 62.9 64.4 56.3 71.2 59.6
Linq-Embed GQR 62.8 65.4 56.7 67.7 61.2
GQR (Top-$K$ only) 62.4 63.5 57.3 68.6 60.1
Qwen3-Embedding GQR 61.0 61.9 54.3 70.2 57.5
GQR (Top-$K$ only) 61.5 63.4 54.3 71.6 56.7
Jina-Embeddings Jina-Embeddings GQR 60.7 61.7 55.3 66.9 58.8
GQR (Top-$K$ only) 59.7 61.7 55.1 65.0 56.8
Linq-Embed GQR 61.2 64.7 57.2 65.7 57.1
GQR (Top-$K$ only) 61.0 64.7 56.9 66.2 56.2
Qwen3-Embedding GQR 59.8 63.2 53.6 67.8 54.4
GQR (Top-$K$ only) 59.0 63.5 53.3 65.0 54.0
Llama-Nemoretriever Jina-Embeddings GQR 64.2 64.5 57.6 74.2 60.4
GQR (Top-$K$ only) 64.2 64.4 57.9 74.1 60.3
Linq-Embed GQR 65.1 66.4 56.8 74.6 62.8
GQR (Top-$K$ only) 64.7 65.4 57.7 74.8 60.8
Qwen3-Embedding GQR 63.3 65.0 55.4 74.1 58.7
GQR (Top-$K$ only) 63.6 65.2 56.9 74.1 58.2
Table 18: Effect of loss function on GQR $\mathrm{NDCG}@5$ performance, over ViDoRe 2.
Avg Biomed Lectures Economics ESG Human ESG Full
Model Complementary Model Loss Variant
Colnomic-Embed Jina-Embeddings Jensen–Shannon 62.7 64.6 56.6 69.5 60.2
Kullback–Leibler (Consensus) 63.0 64.7 57.0 70.3 60.2
Kullback–Leibler (Target) 62.1 64.6 53.3 70.6 60.1
Linq-Embed Jensen–Shannon 63.3 65.3 54.9 71.3 61.9
Kullback–Leibler (Consensus) 62.8 65.4 56.7 67.7 61.2
Kullback–Leibler (Target) 63.8 64.9 57.3 71.3 61.7
Qwen3-Embedding Jensen–Shannon 61.4 63.6 54.3 70.5 57.1
Kullback–Leibler (Consensus) 61.0 61.9 54.3 70.2 57.5
Kullback–Leibler (Target) 61.3 64.2 54.3 70.1 56.7
Jina-Embeddings Jina-Embeddings Jensen–Shannon 60.3 61.7 56.1 67.1 56.5
Kullback–Leibler (Consensus) 60.7 61.7 55.3 66.9 58.8
Kullback–Leibler (Target) 60.9 61.7 55.3 68.5 57.9
Linq-Embed Jensen–Shannon 62.5 63.7 58.7 69.8 57.8
Kullback–Leibler (Consensus) 61.2 64.7 57.2 65.7 57.1
Kullback–Leibler (Target) 61.5 64.7 55.5 68.9 57.1
Qwen3-Embedding Jensen–Shannon 59.0 62.9 53.5 64.7 54.7
Kullback–Leibler (Consensus) 59.8 63.2 53.6 67.8 54.4
Kullback–Leibler (Target) 59.5 63.5 51.9 67.4 55.2
Llama-Nemoretriever Jina-Embeddings Jensen–Shannon 64.0 64.1 57.3 74.3 60.4
Kullback–Leibler (Consensus) 64.2 64.5 57.6 74.2 60.4
Kullback–Leibler (Target) 64.2 64.6 57.2 74.3 60.6
Linq-Embed Jensen–Shannon 65.0 66.0 56.3 74.6 63.1
Kullback–Leibler (Consensus) 65.1 66.4 56.8 74.6 62.8
Kullback–Leibler (Target) 64.7 66.2 56.7 74.3 61.5
Qwen3-Embedding Jensen–Shannon 63.4 65.1 55.7 74.1 58.6
Kullback–Leibler (Consensus) 63.3 65.0 55.4 74.1 58.7
Kullback–Leibler (Target) 63.3 65.1 55.4 74.1 58.7
Table 19: Swapping primary and complementary roles in GQR across model pairs. The first two columns specify the role of each encoder. For each setting we report the absolute score on ViDoRe 2 and the absolute gain relative to the primary encoder alone.
Primary model Complementary model NDCG@5 Gain
Colnomic-7B Jina (text) 63.05 2.8
Jina (text) Colnomic-7B 62.22 8.82
Colnomic-7B Linq-Embed 62.75 2.5
Linq-Embed Colnomic-7B 61.3 6
Colnomic-7B Qwen 3 60.97 0.7
Qwen3 Colnomic-7B 54.4 7.6
Jina (vision) Jina (text) 60.67 3.5
Jina (text) Jina (vision) 59.25 5.85
Jina (vision) Linq-Embed 61.17 4.0
Linq-Embed Jina (vision) 61.37 6.07
Jina (vision) Qwen3 61.17 2.6
Qwen3 Jina (vision) 52.05 5.25
Llama-Nemo Jina (text) 64.17 1.2
Jina (text) Llama-Nemo 59.3 5.9
Llama-Nemo Linq-Embed 65.15 2.2
Linq-Embed Llama-Nemo 60.27 4.97
Llama-Nemo Qwen3 63.3 0.3
Qwen3 Llama-Nemo 52.8 6
Table 20: Tuned values for the GQR step size parameter $\alpha$ (learning rate) over ViDoRe 2.
Selected $\alpha$ value
Primary model Complementary model Biomedical Economics ESG ESG Human
Colnomic-Embed-Multimodal-7B Jina-Embeddings-V4 - Text $5\times 10^{-4}$ $5\times 10^{-3}$ $1\times 10^{-3}$ $5\times 10^{-4}$
Linq-Embed-Mistral $5\times 10^{-4}$ $5\times 10^{-3}$ $1\times 10^{-3}$ $5\times 10^{-3}$
Qwen3-Embedding-4B $\times 10^{-3} 1×1$$\times 10^{-5}$ $\times 10^{-3} 1×1$$\times 10^{-4}$
Jina-Embeddings-V4 Jina-Embeddings-V4 - Text $\times 10^{-5} 1×1$$\times 10^{-3}$ $1\times 10^{-3}$ $5\times 10^{-4}$
Linq-Embed-Mistral $5\times 10^{-4}$ $\times 10^{-3} 1×1$$\times 10^{-4}$ $1\times 10^{-3}$
Qwen3-Embedding-4B $5\times 10^{-4}$ $1\times 10^{-5}$ $5\times 10^{-5}$ $1\times 10^{-3}$
Llama-Nemoretriever-Colembed-3B-V1 Jina-Embeddings-V4 - Text $5\times 10^{-5}$ $\times 10^{-5} 1×1$$\times 10^{-5}$ $1\times 10^{-5}$
Linq-Embed-Mistral $1\times 10^{-4}$ $5\times 10^{-5}$ $\times 10^{-4} 1×1$$\times 10^{-5}$
Qwen3-Embedding-4B $1\times 10^{-4}$ $5\times 10^{-5}$ $\times 10^{-5} 1×1$$\times 10^{-5}$
Table 21: Tuned values for the number of GQR optimization steps parameter $T$ over ViDoRe 2.
Selected $T$ value
Primary model Complementary model Biomedical Economics ESG ESG Human
Colnomic-Embed-Multimodal-7B Jina-Embeddings-V4 - Text 10 25 25 25
Linq-Embed-Mistral 25 25 50 10
Qwen3-Embedding-4B 25 10 10 50
Jina-Embeddings-V4 Jina-Embeddings-V4 - Text 10 50 25 25
Linq-Embed-Mistral 25 25 50 25
Qwen3-Embedding-4B 10 50 50 10
Llama-Nemoretriever-Colembed-3B-V1 Jina-Embeddings-V4 - Text 25 50 50 10
Linq-Embed-Mistral 10 50 10 10
Qwen3-Embedding-4B 10 10 50 10
Table 22: Tuned values for the weight parameter $\alpha$ of the Average Ranking baseline over ViDoRe 2.
Selected $\alpha$ value
Primary model Complementary model Biomedical Economics ESG ESG Human
Colnomic-Embed-Multimodal-7B Jina-Embeddings-V4 - Text 0.8 0.7 0.6 0.8
Linq-Embed-Mistral 0.4 0.5 0.7 0.8
Qwen3-Embedding-4B 0.9 0.9 0.9 0.8
Jina-Embeddings-V4 Jina-Embeddings-V4 - Text 0.9 0.5 0.5 0.8
Linq-Embed-Mistral 0.7 0.3 0.9 0.2
Qwen3-Embedding-4B 0.9 0.9 0.8 0.5
Llama-Nemoretriever-Colembed-3B-V1 Jina-Embeddings-V4 - Text 0.8 0.9 0.9 0.9
Linq-Embed-Mistral 0.9 0.5 0.9 0.9
Qwen3-Embedding-4B 0.9 0.9 0.8 0.9
Table 23: Tuned values for the weight parameter $\alpha$ of the RRF baseline over ViDoRe 2.
Selected $\alpha$ value
Primary model Complementary model Biomedical Economics ESG ESG Human
Colnomic-Embed-Multimodal-7B Jina-Embeddings-V4 - Text 0.9 0.7 0.6 0.9
Linq-Embed-Mistral 0.5 0.5 0.9 0.7
Qwen3-Embedding-4B 0.9 0.9 0.9 0.6
Jina-Embeddings-V4 Jina-Embeddings-V4 - Text 0.9 0.5 0.5 0.8
Linq-Embed-Mistral 0.7 0.3 0.8 0.6
Qwen3-Embedding-4B 0.9 0.9 0.8 0.8
Llama-Nemoretriever-Colembed-3B-V1 Jina-Embeddings-V4 - Text 0.8 0.9 0.9 0.9
Linq-Embed-Mistral 0.9 0.4 0.9 0.9
Qwen3-Embedding-4B 0.9 0.9 0.9 0.9
Table 24: Tuned values for the weight parameter $\alpha$ of the Score Aggregation (Min-Max) baseline over ViDoRe 2.
Selected $\alpha$ value
Primary model Complementary model Biomedical Economics ESG ESG Human
Colnomic-Embed-Multimodal-7B Jina-Embeddings-V4 - Text 0.8 0.4 0.7 0.8
Linq-Embed-Mistral 0.6 0.5 0.6 0.8
Qwen3-Embedding-4B 0.7 0.8 0.9 0.7
Jina-Embeddings-V4 Jina-Embeddings-V4 - Text 0.8 0.7 0.6 0.8
Linq-Embed-Mistral 0.7 0.5 0.8 0.6
Qwen3-Embedding-4B 0.8 0.9 0.8 0.6
Llama-Nemoretriever-Colembed-3B-V1 Jina-Embeddings-V4 - Text 0.8 0.7 0.8 0.9
Linq-Embed-Mistral 0.7 0.8 0.7 0.9
Qwen3-Embedding-4B 0.8 0.9 0.9 0.9
Table 25: Tuned values for the weight parameter $\alpha$ of the Score Aggregation (Softmax) baseline over ViDoRe 2.
Selected $\alpha$ value
Primary model Complementary model Biomedical Economics ESG ESG Human
Colnomic-Embed-Multimodal-7B Jina-Embeddings-V4 - Text 0.6 0.1 0.5 0.7
Linq-Embed-Mistral 0.1 0.2 0.1 0.1
Qwen3-Embedding-4B 0.1 0.2 0.4 0.1
Jina-Embeddings-V4 Jina-Embeddings-V4 - Text 0.9 0.5 0.6 0.6
Linq-Embed-Mistral 0.2 0.1 0.2 0.1
Qwen3-Embedding-4B 0.6 0.5 0.9 0.1
Llama-Nemoretriever-Colembed-3B-V1 Jina-Embeddings-V4 - Text 0.8 0.8 0.9 0.9
Linq-Embed-Mistral 0.1 0.1 0.1 0.9
Qwen3-Embedding-4B 0.2 0.6 0.8 0.9